# RadarLLM: Adapting Pretrained Large Language Models for Marine Radar Target Detection with Preference-aware Loss

### v1

@@ V Conclusion and Future Work
In this study, we propose RadarLLM, a novel framework that integrates a pre-trained LLM with a lightweight reference model for small marine target detection. Employing LoRA adaptation and preference-aware loss, we selectively optimize different feature patches according to their learning values in a lightweight form, effectively reducing model overfitting in low SCR environments. Extensive experimental results demonstrate that RadarLLM significantly outperforms state-of-the-art baselines across sufficient training data, and limited training data settings in various detection scenarios. These results highlight the considerable potential of LLMs in radar signal processing, especially for challenging small marine target detection. Looking ahead, several promising research directions emerge:
@@
    Future studies should investigate the use of more sophisticated open-source pre-trained LLMs, such as Meta’s LLaMA2 or LLaMA3 [42]. These much larger models may offer enhanced parameter transfer capabilities compared to GPT2.

### v2

@@ V Conclusion and Future Work
In this study, we propose RadarLLM, a novel framework that integrates a pre-trained LLM with a lightweight reference model for small marine target detection. Employing LoRA adaptation and preference-aware loss, we selectively optimize different feature patches according to their online evaluated learning values in a lightweight form, effectively reducing model overfitting issues in low SCR environments. Extensive experimental results demonstrate that RadarLLM significantly outperforms state-of-the-art baselines across sufficient training data and limited training data settings in various detection scenarios. These results highlight the considerable potential of LLMs in radar signal processing, especially for challenging small marine target detection. Looking ahead, several promising research directions emerge:
@@
    Future studies should investigate the use of more sophisticated open source pre-trained LLMs, such as LLaMA2 or LLaMA3 [44]. These much larger models may offer enhanced parameter transfer capabilities compared to GPT2.